the time period of a sound wave from a guitar is 2ms. its frequency will be



1
Expert's answer
2021-10-31T18:09:11-0400

Given:

"T=2\:\rm ms=2*10^{-3}\;\rm s"


The frequency of oscillation

"f=\frac{1}{T}=\frac{1}{2*10^{-3}\;\rm s}=500\:\rm Hz"